in a study of distances traveled by buses before the first major engine filure a sample of 191 buses resulted in mean of 96700 miles and a stndard deviation of 37500 mules.at the 0.05 level of significance test the manufcturer's claim that the mean distance traveled before major engine failure is more than 90000 miles.


1
Expert's answer
2021-03-30T10:42:17-0400

We have that:

"n = 191"

"\bar x = 96700"

"s = 37500"

"\alpha = 0.05"

"\mu=90000"


"H_0:\mu = 90000"

"H_a:\mu>90000"


The hypothesis test is right-tailed.

Since the population standard deviation is unknown we use the t-test.

The critical value at the 5% significance level is and 190 df is 1.65

(degrees of freedom df = n – 1 = 191 – 1 = 190)

The critical region is t > 1.65

Test statistic:


"t=\frac{\bar x-\mu}{\frac{s}{\sqrt n}}=\frac{96700-90000}{\frac{37500}{\sqrt 191}}=2.47"

Since 2.47 > 1.65 thus t falls into rejection region therefore we reject the null hypothesis.

At the 5% significance level the data do provide sufficient evidence to confirm the manufacturer's claim. We are 95% confident to conclude that the mean distance traveled before a major engine failure is more than 90,000 miles.